# Mixed FEM for poroelasticity with stress-dependent permeabilityKhan, Lamichhane, Ruiz-Baier & Villa-Fuentes

### v1

@@ 6.1 Reliability of the a posteriori error estimator
$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{\mathbf{div};\Omega}+\|\boldsymbol{u}-\boldsymbol{u}_{h}\|_{0,\Omega}+\|\boldsymbol{\rho}-\boldsymbol{\rho}_{h}\|_{0,\Omega}\leq C_{1}\big{(}\|\mathcal{R}_{1}\|+\|\boldsymbol{f}+\mathbf{div}(\boldsymbol{\sigma}_{h})\|_{0,\Omega}+\|\boldsymbol{\sigma}_{h}-\boldsymbol{\sigma}_{h}^{t}\|_{0,\Omega}\big{)},$
@@
(6.3) $\mathcal{R}_{1}(\boldsymbol{\tau}):=a(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},\boldsymbol{\tau})+b(\boldsymbol{\tau},(\boldsymbol{u}-\boldsymbol{u}_{h},\boldsymbol{\rho}-\boldsymbol{\rho}_{h})),$
with $\mathcal{R}_{1}(\boldsymbol{\tau}_{h})=-c(\boldsymbol{\tau}_{h},p-p_{h})$ for all $\boldsymbol{\tau}_{h}\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}$, and $||\mathcal{R}_{1}||=\sup_{\boldsymbol{0}\neq\boldsymbol{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)}\frac{\mathcal{R}_{1}({\boldsymbol{\tau}})}{\|\boldsymbol{\tau}\|_{\mathbf{div};\Omega}}.$
Proof 6.2.
@@
$\|{\tt e}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|{\tt e}_{\boldsymbol{u}}\|_{0,\Omega}+\|{\tt e}_{\boldsymbol{\rho}}\|_{0,\Omega}$ $\leq C_{1}\sup_{\begin{subarray}{c}\boldsymbol{0}\neq(\boldsymbol{\tau},\boldsymbol{v},{\boldsymbol{\eta}})\\ \quad\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)\end{subarray}}\frac{a({\tt e}_{\boldsymbol{\sigma}},\boldsymbol{\tau})+b(\boldsymbol{\tau},({\tt e}_{\boldsymbol{u}},{\tt e}_{\boldsymbol{\rho}}))+b({\tt e}_{\boldsymbol{\sigma}},(\boldsymbol{v},{\boldsymbol{\eta}}))}{\|\boldsymbol{\tau}\|_{\mathbf{div};\Omega}+\|\boldsymbol{v}\|_{0,\Omega}+\|{\boldsymbol{\eta}}\|_{0,\Omega}}$
$\leq C_{1}\;\Big{(}\sup_{\boldsymbol{0}\neq\boldsymbol{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)}\frac{\mathcal{R}_{1}({\boldsymbol{\tau}})}{\|\boldsymbol{\tau}\|_{\mathbf{div};\Omega}}+\;\sup_{\boldsymbol{0}\neq(\boldsymbol{v},{\boldsymbol{\eta}})\in\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)}\frac{b({\tt e}_{\boldsymbol{\sigma}},(\boldsymbol{v},{\boldsymbol{\eta}}))}{\|\boldsymbol{v}\|_{0,\Omega}+\|{\boldsymbol{\eta}}\|_{0,\Omega}}\Big{)}.$
Then, recalling the definitions of the bilinear form $b$ (cf. (2.1)), using the equation (2.2b), along with the fact that $\int_{\Omega}\boldsymbol{\sigma}_{h}:{\boldsymbol{\eta}}=\frac{1}{2}\int_{\Omega}(\boldsymbol{\sigma}_{h}-\boldsymbol{\sigma}_{h}^{\tt t}):{\boldsymbol{\eta}}$ for ${\boldsymbol{\eta}}\in\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)$, the following estimate holds
$|b({\tt e}_{\boldsymbol{\sigma}},(\boldsymbol{v},{\boldsymbol{\eta}}))|\leq(||\boldsymbol{f}+\mathbf{div}\,\boldsymbol{\sigma}_{h}||_{0,\Omega}+||\boldsymbol{\sigma}_{h}-\boldsymbol{\sigma}_{h}^{\tt t}||_{0,\Omega})(\|\boldsymbol{v}\|_{0,\Omega}+\|{\boldsymbol{\eta}}\|_{0,\Omega}),$
and this gives the asserted inequality.
@@
$\|{\boldsymbol{\varphi}}-{\boldsymbol{\varphi}}_{h}\|_{\mathop{\rm div}\nolimits;\Omega}+\|p-p_{h}\|_{0,\Omega}\leq C_{2}\big{(}\|\mathcal{R}_{2}\|+\|g-c_{0}p_{h}-\dfrac{\alpha}{d\lambda+2\mu}\mathop{\rm tr}\boldsymbol{\sigma}_{h}-\dfrac{d\alpha^{2}}{d\lambda+2\mu}p_{h}+\mathop{\rm div}\nolimits{\boldsymbol{\varphi}}_{h}\|_{0,\Omega}+\gamma\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{\mathbf{div};\Omega}\big{)},$
@@
(6.4) $\mathcal{R}_{2}({\boldsymbol{\psi}}):=\widetilde{a}_{\boldsymbol{\sigma},p}({\boldsymbol{\varphi}}-{\boldsymbol{\varphi}}_{h},{\boldsymbol{\psi}})+\widetilde{b}({\boldsymbol{\psi}},p-p_{h}),$
satisfies $\mathcal{R}_{2}({\boldsymbol{\psi}}_{h})=0$ for all ${\boldsymbol{\psi}}_{h}\in\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}$, and $||\mathcal{R}_{2}||=\sup_{\boldsymbol{0}\neq{\boldsymbol{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)}\frac{\mathcal{R}_{2}({\boldsymbol{\psi}})}{\|{\boldsymbol{\psi}}\|_{\mathop{\rm div}\nolimits;\Omega}}.$
Proof 6.4.
Similarly to Lemma 5.1, using the properties of bilinear forms $\widetilde{a}_{\boldsymbol{\sigma},p}$ and $\widetilde{b}$ (as outlined in equations (3.2b), (2.8b) and (2.9)), along with the insight from [26, Th. 3.4], we establish that there exists $C_{2}>0$ depending on $\kappa_{1}$, $\kappa_{2}$, $C_{\widetilde{a}}$, $\gamma$, $\widetilde{\gamma}$, $\widetilde{\beta}$ such that
$\|{\tt e}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|{\tt e}_{p}\|_{0,\Omega}$ $\leq C_{2}\sup_{\boldsymbol{0}\neq({\boldsymbol{\psi}},q)\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{L}^{2}(\Omega)}\frac{\widetilde{a}_{\boldsymbol{\sigma},p}({\tt e}_{{\boldsymbol{\varphi}}},{\boldsymbol{\psi}})+\widetilde{b}({\boldsymbol{\psi}},{\tt e}_{p})+\widetilde{b}({\tt e}_{{\boldsymbol{\varphi}}},q)-\widetilde{c}({\tt e}_{p},q)}{\|{\boldsymbol{\psi}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|q\|_{0,\Omega}}$
$\leq C_{2}\;\Big{(}\sup_{\boldsymbol{0}\neq{\boldsymbol{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)}\frac{\mathcal{R}_{2}({\boldsymbol{\psi}})}{\|{\boldsymbol{\psi}}\|_{\mathop{\rm div}\nolimits;\Omega}}+\;\sup_{0\neq q\in\mathrm{L}^{2}(\Omega)}\frac{\widetilde{b}({\tt e}_{{\boldsymbol{\varphi}}},q)-\widetilde{c}({\tt e}_{p},q)}{\|q\|_{0,\Omega}}\Big{)}.$
Hence, recalling the definitions of $\widetilde{b},\,\widetilde{c}$, adding ±$\pm\,c\,(\boldsymbol{\sigma}_{h},p)$, and using (2.2d), we arrive at
$|\widetilde{b}({\tt e}_{{\boldsymbol{\varphi}}},q)-\widetilde{c}({\tt e}_{p},q)|\leq\|g-c_{0}p_{h}-\dfrac{\alpha}{d\lambda+2\mu}\mathop{\rm tr}\boldsymbol{\sigma}_{h}-\dfrac{d\alpha^{2}}{d\lambda+2\mu}p_{h}+\mathop{\rm div}\nolimits{\boldsymbol{\varphi}}_{h}\|_{0,\Omega}\|q\|_{0,\Omega}+\gamma\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{\mathbf{div};\Omega}\|q\|_{0,\Omega},$
and therefore, we obtain the desired result.
@@
[…]
@@
$\begin{array}[]{ll}|\mathcal{R}_{1}(\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}}))|\,\leq C_{3}\,\gamma\|p-p_{h}\|_{0,\Omega}\|\mbox{\boldmath{$\xi$}}\|_{1,\Omega}\\ \qquad\displaystyle+C_{3}\biggr{(}\sum_{K\in\mathcal{T}_{h}}h_{K}^{2}||\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}-\boldsymbol{\nabla}\boldsymbol{u}_{h}+\boldsymbol{\rho}_{h}||_{0,K}^{2}+\sum_{E\in\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}h_{E}||\boldsymbol{u}_{\mathrm{D}}-\boldsymbol{u}_{h}||_{0,E}^{2}\biggr{)}^{1/2}\|\mbox{\boldmath{$\xi$}}\|_{1,\Omega}.\end{array}$
Proof 6.6.
@@
From the definition of $\mathcal{R}_{1}$ (cf. (6.3)), adding ±$\pm\,c\,(\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}}),p_{h})$, and using equation (2.2a), we have
@@
$\begin{array}[]{ll}\displaystyle\mathcal{R}_{1}(\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}}))=H(\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}}))-c(\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}}),p)-a(\boldsymbol{\sigma}_{h},\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}}))-b(\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}}),(\boldsymbol{u}_{h},\boldsymbol{\rho}_{h}))\\ \displaystyle\qquad=\langle(\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}}))\mbox{\boldmath{$n$}},\boldsymbol{u}_{\mathrm{D}}\rangle_{\Gamma_{\mathrm{D}}}-\dfrac{\alpha}{d\lambda+2\mu}\int_{\Omega}(p-p_{h})\,\mathop{\rm tr}(\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}}))-\dfrac{\alpha}{d\lambda+2\mu}\int_{\Omega}p_{h}\,\mathop{\rm tr}(\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}}))\\ \displaystyle\qquad-\int_{\Omega}\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}:(\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}}))-\int_{\Omega}\boldsymbol{\rho}_{h}:(\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}}))-\int_{\Omega}\boldsymbol{u}_{h}\cdot\mathbf{div}(\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}})),\end{array}$
[…]
@@
$\begin{array}[]{ll}\displaystyle\mathcal{R}_{1}(\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}}))=\sum_{K\in\mathcal{T}_{h}}\int_{K}(-\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}-\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}+\boldsymbol{\nabla}\boldsymbol{u}_{h}-\boldsymbol{\rho}_{h}):(\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}}))\\ \qquad\displaystyle+\sum_{E\in\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}\langle(\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}}))\mbox{\boldmath{$n$}},\boldsymbol{u}_{\mathrm{D}}-\boldsymbol{u}_{h}\rangle_{E}-\dfrac{\alpha}{d\lambda+2\mu}\int_{\Omega}(p-p_{h})\,\mathop{\rm tr}(\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}}))\\ \qquad\displaystyle\leq\sum_{K\in\mathcal{T}_{h}}\|\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}-\boldsymbol{\nabla}\boldsymbol{u}_{h}+\boldsymbol{\rho}_{h}\|_{0,K}\|\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}})\|_{0,K}\\ \qquad\displaystyle+\sum_{E\in\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}\|\boldsymbol{u}_{\mathrm{D}}-\boldsymbol{u}_{h}\|_{0,E}\|\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}})\|_{0,E}+\gamma\|p-p_{h}\|_{0,\Omega}\|\mbox{\boldmath{$\xi$}}-\mbox{\boldmath{$\Pi$}}_{h}(\mbox{\boldmath{$\xi$}})\|_{0,\Omega},\end{array}$
@@
Let ${\boldsymbol{\chi}}\in\mathbf{H}^{1}_{\Gamma_{\mathrm{N}}}(\Omega):=\{\boldsymbol{w}\in\mathbf{H}^{1}(\Omega):\,\boldsymbol{w}=\boldsymbol{0}\ \text{on}\ \Gamma_{\mathrm{N}}\}$ and assume that $\boldsymbol{u}_{\mathrm{D}}\in\mathbf{H}^{1}(\Gamma_{\mathrm{D}})$. Then, there exists $C_{4}>0$, independent of $h$, such that

### v2

@@ 6.1 Reliability of the a posteriori error estimator
$\|\bm{\sigma}-\bm{\sigma}_{h}\|_{\mathbf{div};\Omega}+\|{\bm{u}}-{\bm{u}}_{h}\|_{0,\Omega}+\|\bm{\rho}-\bm{\rho}_{h}\|_{0,\Omega}\leq C_{1}\big{(}\|\mathcal{R}_{1}\|+\|\bm{f}+\mathbf{div}(\bm{\sigma}_{h})\|_{0,\Omega}+\|\bm{\sigma}_{h}-\bm{\sigma}_{h}^{t}\|_{0,\Omega}\big{)},$
@@
$\mathcal{R}_{1}(\bm{\tau}):=a(\bm{\sigma}-\bm{\sigma}_{h},\bm{\tau})+b(\bm{\tau},({\bm{u}}-{\bm{u}}_{h},\bm{\rho}-\bm{\rho}_{h})),$ (6.3)
with $\mathcal{R}_{1}(\bm{\tau}_{h})=-c(\bm{\tau}_{h},p-p_{h})$ for all $\bm{\tau}_{h}\in\mathbb{H}^{\bm{\sigma}}_{h}$, and $||\mathcal{R}_{1}||=\sup_{\bm{0}\neq\bm{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)}\frac{\mathcal{R}_{1}({\bm{\tau}})}{\|\bm{\tau}\|_{\mathbf{div};\Omega}}.$
Proof.
@@
$\|{\tt e}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|{\tt e}_{{\bm{u}}}\|_{0,\Omega}+\|{\tt e}_{\bm{\rho}}\|_{0,\Omega}\leq C_{1}\!\!\!\!\!\!\!\!\sup_{\begin{subarray}{c}\bm{0}\neq(\bm{\tau},\mathbf{v},{\bm{\eta}})\\ \quad\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)\end{subarray}}\!\!\!\!\frac{a({\tt e}_{\bm{\sigma}},\bm{\tau})+b(\bm{\tau},({\tt e}_{{\bm{u}}},{\tt e}_{\bm{\rho}}))+b({\tt e}_{\bm{\sigma}},(\mathbf{v},{\bm{\eta}}))}{\|\bm{\tau}\|_{\mathbf{div};\Omega}+\|\mathbf{v}\|_{0,\Omega}+\|{\bm{\eta}}\|_{0,\Omega}}$
$\qquad\leq C_{1}\;\Big{(}\sup_{\bm{0}\neq\bm{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)}\frac{\mathcal{R}_{1}({\bm{\tau}})}{\|\bm{\tau}\|_{\mathbf{div};\Omega}}+\;\sup_{\bm{0}\neq(\mathbf{v},{\bm{\eta}})\in\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)}\frac{b({\tt e}_{\bm{\sigma}},(\mathbf{v},{\bm{\eta}}))}{\|\mathbf{v}\|_{0,\Omega}+\|{\bm{\eta}}\|_{0,\Omega}}\Big{)}.$
Then, recalling the definitions of the bilinear form $b$ (cf. (2.1)), using the equation (2.2b), along with the fact that $\int_{\Omega}\bm{\sigma}_{h}:{\bm{\eta}}=\frac{1}{2}\int_{\Omega}(\bm{\sigma}_{h}-\bm{\sigma}_{h}^{\tt t}):{\bm{\eta}}$ for ${\bm{\eta}}\in\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)$, the following estimate holds
$|b({\tt e}_{\bm{\sigma}},(\mathbf{v},{\bm{\eta}}))|\leq(||\bm{f}+\mathbf{div}\,\bm{\sigma}_{h}||_{0,\Omega}+||\bm{\sigma}_{h}-\bm{\sigma}_{h}^{\tt t}||_{0,\Omega})(\|\mathbf{v}\|_{0,\Omega}+\|{\bm{\eta}}\|_{0,\Omega}),$
and this gives the asserted inequality. ∎
@@
$\|{\bm{\varphi}}-{\bm{\varphi}}_{h}\|_{\mathop{\rm div}\nolimits;\Omega}+\|p-p_{h}\|_{0,\Omega}$
$\quad\leq C_{2}\big{(}\|\mathcal{R}_{2}\|+\|g-c_{0}p_{h}-\dfrac{\alpha}{d\lambda+2\mu}\mathop{\rm tr}\bm{\sigma}_{h}-\dfrac{d\alpha^{2}}{d\lambda+2\mu}p_{h}+\mathop{\rm div}\nolimits{\bm{\varphi}}_{h}\|_{0,\Omega}+\gamma\|\bm{\sigma}-\bm{\sigma}_{h}\|_{\mathbf{div};\Omega}\big{)},$
@@
$\mathcal{R}_{2}({\bm{\psi}}):=\widetilde{a}_{\bm{\sigma},p}({\bm{\varphi}}-{\bm{\varphi}}_{h},{\bm{\psi}})+\widetilde{b}({\bm{\psi}},p-p_{h}),$
satisfies $\mathcal{R}_{2}({\bm{\psi}}_{h})=0$ for all ${\bm{\psi}}_{h}\in\mathbf{H}^{{\bm{\varphi}}}_{h}$, and $||\mathcal{R}_{2}||=\sup_{\bm{0}\neq{\bm{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)}\frac{\mathcal{R}_{2}({\bm{\psi}})}{\|{\bm{\psi}}\|_{\mathop{\rm div}\nolimits;\Omega}}.$
Proof.
Similarly to Lemma 5.1, using the properties of bilinear forms $\widetilde{a}_{\bm{\sigma},p}$ and $\widetilde{b}$ (as outlined in equations (3.2b), (2.8b) and (2.9)), along with the insight from [26, Th. 3.4], we establish that there exists $C_{2}>0$ depending on $\kappa_{1}$, $\kappa_{2}$, $C_{\widetilde{a}}$, $\gamma$, $\widetilde{\gamma}$, $\widetilde{\beta}$ such that
$\|{\tt e}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|{\tt e}_{p}\|_{0,\Omega}$ $\leq C_{2}\sup_{\bm{0}\neq({\bm{\psi}},q)\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{L}^{2}(\Omega)}\frac{\widetilde{a}_{\bm{\sigma},p}({\tt e}_{{\bm{\varphi}}},{\bm{\psi}})+\widetilde{b}({\bm{\psi}},{\tt e}_{p})+\widetilde{b}({\tt e}_{{\bm{\varphi}}},q)-\widetilde{c}({\tt e}_{p},q)}{\|{\bm{\psi}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|q\|_{0,\Omega}}$
$\leq C_{2}\;\Big{(}\sup_{\bm{0}\neq{\bm{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)}\frac{\mathcal{R}_{2}({\bm{\psi}})}{\|{\bm{\psi}}\|_{\mathop{\rm div}\nolimits;\Omega}}+\;\sup_{0\neq q\in\mathrm{L}^{2}(\Omega)}\frac{\widetilde{b}({\tt e}_{{\bm{\varphi}}},q)-\widetilde{c}({\tt e}_{p},q)}{\|q\|_{0,\Omega}}\Big{)}.$
Hence, recalling the definitions of $\widetilde{b},\,\widetilde{c}$, adding ±$\pm\,c\,(\bm{\sigma}_{h},p)$, and using (2.2d), we arrive at
$|\widetilde{b}({\tt e}_{{\bm{\varphi}}},q)-\widetilde{c}({\tt e}_{p},q)|\leq\|g-c_{0}p_{h}-\dfrac{\alpha}{d\lambda+2\mu}\mathop{\rm tr}\bm{\sigma}_{h}-\dfrac{d\alpha^{2}}{d\lambda+2\mu}p_{h}+\mathop{\rm div}\nolimits{\bm{\varphi}}_{h}\|_{0,\Omega}\|q\|_{0,\Omega}+\gamma\|\bm{\sigma}-\bm{\sigma}_{h}\|_{\mathbf{div};\Omega}\|q\|_{0,\Omega},$
and therefore, we obtain the desired result. ∎
@@
There exists a positive constant $C_{3}$, independent of $h$, such that for each $\bm{\xi}\in\mathbb{H}^{1}(\Omega)$ there holds
@@
$\begin{array}[]{ll}|\mathcal{R}_{1}(\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi}))|\,\leq C_{3}\,\gamma\|p-p_{h}\|_{0,\Omega}\|\bm{\xi}\|_{1,\Omega}\\ \quad\displaystyle+C_{3}\biggr{(}\sum_{K\in\mathcal{T}_{h}}h_{K}^{2}||\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}-\bm{\nabla}{\bm{u}}_{h}+\bm{\rho}_{h}||_{0,K}^{2}+\sum_{E\in\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}h_{E}||{\bm{u}}_{\mathrm{D}}-{\bm{u}}_{h}||_{0,E}^{2}\biggr{)}^{1/2}\|\bm{\xi}\|_{1,\Omega}.\end{array}$
Proof.
@@
From the definition of $\mathcal{R}_{1}$ (cf. (6.3)), adding ±$\pm\,c\,(\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi}),p_{h})$, and using equation (2.2a), we have
@@
$\begin{array}[]{ll}\displaystyle\mathcal{R}_{1}(\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi}))=H(\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi}))-c(\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi}),p)-a(\bm{\sigma}_{h},\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi}))-b(\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi}),({\bm{u}}_{h},\bm{\rho}_{h}))\\ \displaystyle\quad=\langle(\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi}))\mbox{\boldmath{$n$}},{\bm{u}}_{\mathrm{D}}\rangle_{\Gamma_{\mathrm{D}}}-\dfrac{\alpha}{d\lambda+2\mu}\int_{\Omega}(p-p_{h})\,\mathop{\rm tr}(\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi}))-\dfrac{\alpha}{d\lambda+2\mu}\int_{\Omega}p_{h}\,\mathop{\rm tr}(\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi}))\\ \displaystyle\qquad-\int_{\Omega}\mathcal{C}^{-1}\bm{\sigma}_{h}:(\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi}))-\int_{\Omega}\bm{\rho}_{h}:(\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi}))-\int_{\Omega}{\bm{u}}_{h}\cdot\mathbf{div}(\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi})),\end{array}$
then, applying a local integration by parts to the last term above, using the identity $\int_{E}\mbox{\boldmath{$\Pi$}}_{h}(\bm{\tau})\mbox{\boldmath{$n$}}\cdot\bm{\xi}=\int_{E}\bm{\tau}\mbox{\boldmath{$n$}}\cdot\bm{\xi}$, for all $\bm{\xi}\in\mathbf{P}_{k}(E)$, for all edge $E$ of $\mathcal{T}_{h}$, the fact that ${\bm{u}}_{\mathrm{D}}\in\mathbf{L}^{2}(\Gamma_{\mathrm{D}})$, and the Cauchy-Schwarz inequality, we obtain
@@
$\begin{array}[]{ll}\displaystyle\mathcal{R}_{1}(\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi}))=\sum_{K\in\mathcal{T}_{h}}\int_{K}(-\mathcal{C}^{-1}\bm{\sigma}_{h}-\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}+\bm{\nabla}{\bm{u}}_{h}-\bm{\rho}_{h}):(\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi}))\\ \qquad\qquad\displaystyle+\sum_{E\in\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}\langle(\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi}))\mbox{\boldmath{$n$}},{\bm{u}}_{\mathrm{D}}-{\bm{u}}_{h}\rangle_{E}-\dfrac{\alpha}{d\lambda+2\mu}\int_{\Omega}(p-p_{h})\,\mathop{\rm tr}(\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi}))\\ \qquad\displaystyle\leq\sum_{K\in\mathcal{T}_{h}}\|\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}-\bm{\nabla}{\bm{u}}_{h}+\bm{\rho}_{h}\|_{0,K}\|\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi})\|_{0,K}\\ \qquad\qquad\displaystyle+\sum_{E\in\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}\|{\bm{u}}_{\mathrm{D}}-{\bm{u}}_{h}\|_{0,E}\|\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi})\|_{0,E}+\gamma\|p-p_{h}\|_{0,\Omega}\|\bm{\xi}-\mbox{\boldmath{$\Pi$}}_{h}(\bm{\xi})\|_{0,\Omega},\end{array}$
@@
Let ${\bm{\chi}}\in\mathbf{H}^{1}_{\Gamma_{\mathrm{N}}}(\Omega):=\{\bm{w}\in\mathbf{H}^{1}(\Omega):\,\bm{w}=\bm{0}\ \text{on}\ \Gamma_{\mathrm{N}}\}$ and assume that ${\bm{u}}_{\mathrm{D}}\in\mathbf{H}^{1}(\Gamma_{\mathrm{D}})$. Then, there exists $C_{4}>0$, independent of $h$, such that